OpenStudy (anonymous):

find the sum of negative 5 and positive 12 and find the sum of negative 8 and negative 11 and find the sum of positive 12 and negative 19

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

-5+12=7 -8-11=-19 12-19=-7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

5+(-2) an -8+0 an -3+(-8) an 18+(-12)